Break-even sales and sales to realize operating income

For the current year ended March 31, Cosgrove Company expects fixed costs of $27,600,000, a unit variable cost of $805, and a unit selling price of $1,150.

a. Compute the anticipated break-even sales (units).
____units

b. Compute the sales (units) required to realize operating income of $5,175,000.
_____units
Answer #1

Contribution margin=Sales-Variable cost

=(1150-805)=$345 per unit

a.Breakeven=Fixed expenses/Contribution margin

=27,600,000/345

=80000 units

b.Target Contribution margin=Fixed cost+Target profits

=27,600,000+5,175,000

=$32775000

Hence unit sales=Target Contribution margin/Contribution margin per unit

=32775000/345

=95000 units